# Multi-Agent Quantum Reinforcement Learning using Evolutionary Optimization

### v2

@@ 4 APPROACH
$\mathcal{P}_{0}\leftarrow$ Initialize population $\eta$ with random $\theta$ for $g\in\{0,1,...,\mu\}$ do
@@
             Reset testing environment Score $S_{t,i}\leftarrow 0$ for $t\in\{0,1,...,\kappa\}$ do
                   Use policy of agent $i$ for all agents in env Select action $a_{t}\leftarrow\text{argmax}\ VQC_{\theta}(s_{t})$ Execute environment step with action $a_{t}$ Observe reward $r_{t}$ and next state $s_{t+1}$ $S_{t,i}\leftarrow S_{t,i}+r_{t}$
@@
      $\lambda\leftarrow$ Select top $\tau$ agents based on $S_{t,i}$ Keep top agent based on $S_{t,i}$ Recombine $\eta-1$ new agents out of $\lambda$ Mutate $\eta-1$ generated agents $\mathcal{P}_{g+1}\leftarrow$ $\eta-1$ generated agents + top agent
@@ 6.1 Comparing Generational Evolution Strategies
Beyond score comparison, we evaluated the average number of coins collected during the experiments. As inferred from the scores, the mutation-only strategy consistently collects more coins, as shown in 4(a). Although there are periods where the strategies yield almost identical coin counts, at other times, a gap of up to $2$ coins is evident. On average, the combined strategies lag slightly behind the mutation-only in terms of coin collection.
The layerwise crossover’s initial surge in Figure 4 correlates with the uptrend in collected coins shown in 4(a) and the coin rate detailed in 4(c). When comparing the strategies based on these metrics, mutation-only consistently achieves the highest coin rate over all generations and collects the most coins, accounting for its superior average reward. The random crossover strategy, while collecting more coins than the layerwise approach, has a significantly reduced coin rate, resulting in diminished overall rewards.
Exploring the coin rate, depicted in 4(c), the layerwise strategy leads until the 90th generation. After that, its rate declines, while the mutation-only strategy exhibits a gradual, consistent rise. A higher coin rate indicates enhanced agent cooperation within the testing environment. This coin rate, combined with the number of coins collected, determines an agent’s reward.
@@ 6.2 Assessing Varying Layer Counts
For a comprehensive understanding, we next probe the average coin collection in 6(a). The 4-layer VQC consistently tops the coin collection metric, progressing from just below $7$ to $8$. The 6-layer VQC commences at $6.5$, dips to $5.2$ by the 23rd generation, and then rises to 8 by the 165th generation. The 8-layer VQC, despite securing the highest average reward, begins at $6$ and only stabilizes around $8$ after the 180th generation. The 16-layer VQC, after an initial dip, witnesses a rapid increase from the 24th to 103rd generation, briefly declines, and then fluctuates around $8$ coins. Towards the concluding generations, VQCs with more than 4 layers converge to collect approximately 8 coins.
Analyzing the own coin count in 6(b), we observe that, except for the 4-layer VQC, all VQCs initially decline before ascending. The 4-layer VQC displays a steady yet modest climb, concluding at a count of $6.5$. The 6-layer VQC takes the longest to commence its ascent, eventually oscillating around a count of $7.2$. The 8-layer VQC initiates its climb earlier, achieving a slightly higher count of $7.5$ by the end. The 16-layer VQC, notable for its rÏapid early ascent, consistently hovers around a count of $7$ after the 100th generation. Among the VQCs, the 4-layer variant lags, collecting over one own coin fewer than its counterparts.
@@ 6.3.1 Comparing VQC and Random
First, we compare the results of our VQC approaches to the results of the random baseline. In Figure 8, the score for the random acting agents is approximately 00, since the cooperative sequential coin game is a zero-sum game. The evolutionary-trained VQC approache, however, perform significantly better leading to an average score around $7$. The total coins collected depicted in 8(a) suggests that in contrast to the random agents, the VQC agents successfully learn to collect coins. The own coins collected correlates with number of collected coins (8(b)). In 8(c) we can see that in neither case the cooperation increases over time. In summary, trained agents performed significantly better than random on all metrics, indicating that the training was successful.
@@ 6.3.2 Comparing VQC and Small NN
As depicted in Figure 8, a better result is achieved by the VQC approach compared to random. On this basis, we compare the performance of this VQC approach to that of a neural network with a comparable number of parameters. Here we exploit the higher expressive power of VQCs compared to conventional neural networks [Chen et al., 2022]. Similar to [Chen et al., 2022], we define the expressive power as the capacity to represent particular functions with a constrained number of parameters. Note that the VQC has $148$ parameters (3 * 6 * 8 + 4). The neural network uses two hidden layers with dimension 3 and 4 respectively, resulting in a parameter count of $147$. Both the neural network and the VQC, are trained with mutation only with mutation power $\sigma=0.01$. In Figure 8, we can see that the neural network reward fluctuates in the range of $2.5$ to $3$. As previously discussed in the last section, the VQC approach exhibits a slow learning curve leading to a significant higher score therefore consequently outperforming this neural network. The inferior performance can be explained by the small number of hidden units and parameters present in neural networks. Typically, the number of hidden units is chosen much higher. Further evidence of the neural network’s deficiency is provided by the average number of coins collected. As shown in 8(a), the NN’s number of collected coins is below the average score of the random agents until generation $115$ and after that slightly over it. In comparison, the VQC with the same number of parameters collects two times as many coins on average. The neural network is able to outperform random agents on the basis of its collected own coins, what can be seen in 8(b), leading to the better performance regarding the own coin rate (8(c)). In terms of collected own coins and the own coin rate, the neural network performs significantly worse than the VQC with nearly the same number of parameters. A neural network with this few hidden units and, consequently, parameters is not able learn in the coin game environment successfully. This demonstrates the power of VQCs for RL archiving significantly higher with the same amount of parameters.
@@ 6.3.3 Comparing VQC and Big NN
In the previous section, we observed that a neural network with the same number of parameters as the VQC of our approach cannot match the VQC’s performance. We will now compare the results with a neural network that has significantly more parameters. Again, mutation only is used for the evolution of subsequent generations in both cases and the mutation power is $\sigma=0.01$. We chose a fully connected NN with two hidden layers of size $64$, resulting in a parameter count of $6788$. If we first examine the reward of the neural network and the VQC, we can see in Figure 8 that both produce very similar results over time. Initially, the VQC has a slightly higher score. From generation $50$ onward, there are only minor differences in terms of average score. Overall, the two strategies yield a score of approximatly $7$. On the basis of our experiments, we can say that the VQC achieves nearly identical performance in the Coin Game environment compared to a neural network that has $46$ times more parameters. In 8(a), the initial value of the VQC method is again higher than that of the neural network. Due to a steeper learning curve, the neural network is able to compensate the lower starting value and achieves only a slightly smaller number of collected coins. From there are no discernible differences between the two models. A similar performance of the two approaches can be seen in 8(b), where the average number of own coins collected is shown. In Terms of own coin rate, at first, the VQC archives a slightly higher score. However, at Generation 25, the neural network initially achieves a better own coin rate before being slightly lower between Generation 80 and 162. In the end, the neural network is slightly better in terms of the own coin rate.
@@ ACKNOWLEDGEMENTS
This work is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.

### v3

@@ 4 APPROACH
$\mathcal{P}_{0}\leftarrow$ Initialize population $\eta$ with random $\theta$
for $g\in\{0,1,...,\mu\}$ do
@@
             Reset testing environment
             Score $S_{t,i}\leftarrow 0$
             for $t\in\{0,1,...,\kappa\}$ do
                   Use policy of agent $i$ for all agents in env
                   Select action $a_{t}\leftarrow\text{argmax}\ VQC_{\theta}(s_{t})$
                   Execute environment step with action $a_{t}$
                   Observe reward $r_{t}$ and next state $s_{t+1}$
                   $S_{t,i}\leftarrow S_{t,i}+r_{t}$
@@
      $\lambda\leftarrow$ Select top $\tau$ agents based on $S_{t,i}$
       Keep top agent based on $S_{t,i}$
       Recombine $\eta-1$ new agents out of $\lambda$
       Mutate $\eta-1$ generated agents
       $\mathcal{P}_{g+1}\leftarrow$ $\eta-1$ generated agents + top agent
@@ 6.1 Comparing Generational Evolution Strategies
Beyond score comparison, we evaluated the average number of coins collected during the experiments. As inferred from the scores, the mutation-only strategy consistently collects more coins, as shown in 5(a). Although there are periods where the strategies yield almost identical coin counts, at other times, a gap of up to $2$ coins is evident. On average, the combined strategies lag slightly behind the mutation-only in terms of coin collection.
The layerwise crossover’s initial surge in Figure 4 correlates with the uptrend in collected coins shown in 5(a) and the coin rate detailed in 5(c). When comparing the strategies based on these metrics, mutation-only consistently achieves the highest coin rate over all generations and collects the most coins, accounting for its superior average reward. The random crossover strategy, while collecting more coins than the layerwise approach, has a significantly reduced coin rate, resulting in diminished overall rewards.
Exploring the coin rate, depicted in 5(c), the layerwise strategy leads until the 90th generation. After that, its rate declines, while the mutation-only strategy exhibits a gradual, consistent rise. A higher coin rate indicates enhanced agent cooperation within the testing environment. This coin rate, combined with the number of coins collected, determines an agent’s reward.
@@ 6.2 Assessing Varying Layer Counts
For a comprehensive understanding, we next probe the average coin collection in 7(a). The 4-layer VQC consistently tops the coin collection metric, progressing from just below $7$ to $8$. The 6-layer VQC commences at $6.5$, dips to $5.2$ by the 23rd generation, and then rises to 8 by the 165th generation. The 8-layer VQC, despite securing the highest average reward, begins at $6$ and only stabilizes around $8$ after the 180th generation. The 16-layer VQC, after an initial dip, witnesses a rapid increase from the 24th to 103rd generation, briefly declines, and then fluctuates around $8$ coins. Towards the concluding generations, VQCs with more than 4 layers converge to collect approximately 8 coins.
Analyzing the own coin count in 7(b), we observe that, except for the 4-layer VQC, all VQCs initially decline before ascending. The 4-layer VQC displays a steady yet modest climb, concluding at a count of $6.5$. The 6-layer VQC takes the longest to commence its ascent, eventually oscillating around a count of $7.2$. The 8-layer VQC initiates its climb earlier, achieving a slightly higher count of $7.5$ by the end. The 16-layer VQC, notable for its rÏapid early ascent, consistently hovers around a count of $7$ after the 100th generation. Among the VQCs, the 4-layer variant lags, collecting over one own coin fewer than its counterparts.
@@ 6.3.1 Comparing VQC and Random
First, we compare the results of our VQC approaches to the results of the random baseline. In Figure 8, the score for the random acting agents is approximately 00, since the cooperative sequential coin game is a zero-sum game. The evolutionary-trained VQC approache, however, perform significantly better leading to an average score around $7$. The total coins collected depicted in 9(a) suggests that in contrast to the random agents, the VQC agents successfully learn to collect coins. The own coins collected correlates with number of collected coins (9(b)). In 9(c) we can see that in neither case the cooperation increases over time. In summary, trained agents performed significantly better than random on all metrics, indicating that the training was successful.
@@ 6.3.2 Comparing VQC and Small NN
As depicted in Figure 8, a better result is achieved by the VQC approach compared to random. On this basis, we compare the performance of this VQC approach to that of a neural network with a comparable number of parameters. Here we exploit the higher expressive power of VQCs compared to conventional neural networks [Chen et al., 2022]. Similar to [Chen et al., 2022], we define the expressive power as the capacity to represent particular functions with a constrained number of parameters. Note that the VQC has $148$ parameters (3 * 6 * 8 + 4). The neural network uses two hidden layers with dimension 3 and 4 respectively, resulting in a parameter count of $147$. Both the neural network and the VQC, are trained with mutation only with mutation power $\sigma=0.01$. In Figure 8, we can see that the neural network reward fluctuates in the range of $2.5$ to $3$. As previously discussed in the last section, the VQC approach exhibits a slow learning curve leading to a significant higher score therefore consequently outperforming this neural network. The inferior performance can be explained by the small number of hidden units and parameters present in neural networks. Typically, the number of hidden units is chosen much higher. Further evidence of the neural network’s deficiency is provided by the average number of coins collected. As shown in 9(a), the NN’s number of collected coins is below the average score of the random agents until generation $115$ and after that slightly over it. In comparison, the VQC with the same number of parameters collects two times as many coins on average. The neural network is able to outperform random agents on the basis of its collected own coins, what can be seen in 9(b), leading to the better performance regarding the own coin rate (9(c)). In terms of collected own coins and the own coin rate, the neural network performs significantly worse than the VQC with nearly the same number of parameters. A neural network with this few hidden units and, consequently, parameters is not able learn in the coin game environment successfully. This demonstrates the power of VQCs for RL archiving significantly higher with the same amount of parameters.
@@ 6.3.3 Comparing VQC and Big NN
In the previous section, we observed that a neural network with the same number of parameters as the VQC of our approach cannot match the VQC’s performance. We will now compare the results with a neural network that has significantly more parameters. Again, mutation only is used for the evolution of subsequent generations in both cases and the mutation power is $\sigma=0.01$. We chose a fully connected NN with two hidden layers of size $64$, resulting in a parameter count of $6788$. If we first examine the reward of the neural network and the VQC, we can see in Figure 8 that both produce very similar results over time. Initially, the VQC has a slightly higher score. From generation $50$ onward, there are only minor differences in terms of average score. Overall, the two strategies yield a score of approximatly $7$. On the basis of our experiments, we can say that the VQC achieves nearly identical performance in the Coin Game environment compared to a neural network that has $46$ times more parameters. In 9(a), the initial value of the VQC method is again higher than that of the neural network. Due to a steeper learning curve, the neural network is able to compensate the lower starting value and achieves only a slightly smaller number of collected coins. From there are no discernible differences between the two models. A similar performance of the two approaches can be seen in 9(b), where the average number of own coins collected is shown. In Terms of own coin rate, at first, the VQC archives a slightly higher score. However, at Generation 25, the neural network initially achieves a better own coin rate before being slightly lower between Generation 80 and 162. In the end, the neural network is slightly better in terms of the own coin rate.
@@ ACKNOWLEDGEMENTS
This work is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. This publication was created as part of the Q-Grid project (13N16179) under the “quantum technologies - from basic research to market” funding program, supported by the German Federal Ministry of Education and Research.